# Why These Documents? Explainable Generative Retrieval with Hierarchical Category Paths

### v1

@@ Abstract.
Generative retrieval has recently emerged as a new alternative of traditional information retrieval approaches. However, existing generative retrieval methods directly decode docid when a query is given, making it impossible to provide users with explanations as an answer for “Why this document is retrieved?”. To address this limitation, we propose Hierarchical Category Path-Enhanced Generative Retrieval (HyPE), which enhances explainability by generating hierarchical category paths step-by-step before decoding docid. HyPE leverages hierarchical category paths as explanation, progressing from broad to specific semantic categories. This approach enables diverse explanations for the same document depending on the query by using shared category paths between the query and the document, and provides reasonable explanation by reflecting the document’s semantic structure through a coarse-to-fine manner. HyPE constructs category paths with external high-quality semantic hierarchy, leverages LLM to select appropriate candidate paths for each document, and optimizes the generative retrieval model with path-augmented dataset. During inference, HyPE utilizes path-aware reranking strategy to aggregate diverse topic information, allowing the most relevant documents to be prioritized in the final ranked list of docids. Our extensive experiments demonstrate that HyPE not only offers a high level of explainability but also improves the retrieval performance in the document retrieval task.
Information Retrieval, Generative Retrieval, Explainability
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and ranking
1. Introduction
@@
Figure 1. Existing generative retrieval methods fail to explain why specific documents are retrieved, as they directly decode docid (Upper). In contrast, our proposed HyPE method provides clear explanations by generating query-related hierarchical category paths leading to the docid (Lower).
Information retrieval (IR) systems play a vital role in helping users navigate the vast amount of online information. A fundamental task of these systems is document retrieval, which focuses on searching for and ranking documents that are relevant to a given query from a large corpus. Over the past few years, document retrieval has evolved rapidly and has become pivotal in several NLP downstream tasks, including web search engines, question answering, and dialogue systems. In the field of document retrieval, sparse retrieval (Robertson and Zaragoza, 2009; Formal et al., 2021) based on term-matching and dense retrieval (Karpukhin et al., 2020; Khattab and Zaharia, 2020; Xiong et al., 2021) based on semantic representations have been widely employed.
@@
With the significant progress of language models, generative retrieval has recently emerged as a new paradigm in document retrieval. Generate retrieval aims to directly generate document identifier (i.e., docid) for a given query by leveraging pre-trained generative models such as BART (Lewis et al., 2020) and T5 (Raffel et al., 2020). This paradigm enables end-to-end optimization of the retrieval process, allowing for fine-grained interaction between the input query and docid, and significantly reduces memory usage by fully leveraging the parametric memory of a single pre-trained generative model.
@@
Despite these advantages, generative retrieval continues to face the challenge of determining how to construct docid that effectively represent documents. As the docid serves as a representation of the entire document, defining one that accurately encapsulates the document’s contents is both crucial and challenging. Existing workds on generative retrieval have categorized docid into two types: semantic docid and lexical docid. Semantic docid represents each document as a unique identifier, either a single number or a series of numbers (e.g., ‘0-5-2’), derived from the dense representations encoded by a pre-trained text encoder (Devlin et al., 2019) and clustering techniques such as hierarchical k-means clustering (Tay et al., 2022; Wang et al., 2022; Sun et al., 2023) and product quantization (Zhou et al., 2022). On the other hand, lexical docid represents each document as human-readable text, such as titles (Cao et al., 2021), keywords (Zhang et al., 2023; Wang et al., 2023; Lee et al., 2023) and pseudo queries (Tang et al., 2023; Li et al., 2023).
@@
Unlike semantic docid, lexical docid achieves interpretability by providing human-readable identifiers, but they still lack in explainability regarding why these documents were retrieved. In Figure 1 upper, existing generative retrieval methods directly decode docid when a query is given, making it impossible to provide users with explanations as an answer for “Why this document is retrieved?”. While users can recognize what a lexical docid represents, it is not always obvious why a certain docid is decoded for a specific query. The lack of explainability in retrieval systems can undermine the reliability of the retrieved documents and make it more challenging for users to explore additional information related to a specific query (Anand et al., 2022), making it a crucial issue to address. To address the aforementioned limitation, our research aims to design a generative retrieval framework that can provide retrieved results with reasonable explanations for the user’s query.
@@
In this work, we propose Hierarchical Category Path-Enhanced Generative Retrieval (HyPE), which enhances explainability by generating hierarchical category paths step-by-step before decoding docid. Motivated by semantic docid, HyPE employs hierarchical category paths as explanation, progressing from broad to specific semantic categories. In Figure 1 lower, when queries about document “Dubai” are given, HyPE uses category paths like “Government ¿ Government by cities” or “Economy ¿ Economy by cities” to explain why document “Dubai” is retrieved for each query. This approach 1) enables diverse explanations for the same document depending on the query by using shared category paths between the query and the document, and 2) provides more reasonable and insightful explanation by reflecting the document’s semantic structure through a coarse-to-fine manner. Additionally, HyPE 3) can employ effective reranking of the retrieved results by leveraging multiple paths, which helps improve retrieval performance.
@@
Specifically, HyPE consists of the following three steps: 1) constructing category paths based on an external high-quality semantic hierarchy and selecting appropriate candidate paths for each document using Large Language Models (LLM), 2) building a path-augmented dataset with candidate paths, and 3) optimizing a model with the path-augmented dataset. During inference phase, HyPE conducts a pseudo-reasoning [^1] Generating the hierarchical path resembles step-by-step reasoning. However, unlike natural language-based reasoning in LLM, we use the term “pseudo-reasoning” because the path structure is more akin to pseudo-code. process by generating the hierarchical category path step-by-step to decode docid, allowing it to serve as an explanation which enhances explainability. Additionally, HyPE employs path-aware reranking strategy, which simultaneously considers multiple pseudo-reasoning paths for each query. This strategy helps build a more robust retrieval system by capturing the semantic information of multiple category paths, thereby improving overall retrieval performance.
@@
Our extensive experiments demonstrate that HyPE not only offers a high level of explainability but also improves the retrieval performance in document retrieval task. Additionally, HyPE can be applied orthogonally to various docid types (e.g., title, keywords), making it a versatile framework that can be seamlessly integrated into different generative retrieval systems. For reproducibility, our codes are publicly available at the anonymous github repository.222 https://github.com/augustinLib/HyPE-Why-These-Documents
The contributions of our work are summarized as follows:
@@
  • We propose a new reranking strategy called path-aware reranking, which considers multiple category paths simultaneously to determine the final ranked list of docids.
@@
2. Related Work
2.1. Document Retrieval
Document retrieval is a representative task in information retrieval (IR) that involves searching for documents relevant to a given query within a document corpus. Traditionally, it has been accomplished through sparse retrieval and dense retrieval. Sparse retrieval (Robertson and Zaragoza, 2009; Formal et al., 2021) employs term-matching scoring methods, such as BM25 (Robertson and Zaragoza, 2009), to calculate relevance scores for retrieving documents. However, these methods struggle with the issue of lexical mismatch (Lin et al., 2020) due to terms such as homonyms and synonyms. To address these limitations, dense retrieval  (Karpukhin et al., 2020; Khattab and Zaharia, 2020; Xiong et al., 2021) has emerged with the recent development of pre-trained language models (PLMs). This approach utilizes encoder-based PLM, such as BERT (Devlin et al., 2019), to obtain dense representations for both queries and documents, and retrieve relevant documents through an Approximate Nearest Neighbor (ANN) index (Malkov and Yashunin, 2016). Although dense retrieval methods have demonstrated impressive performance by capturing deep semantic information that sparse retrieval methods cannot, it still has the limitation of not being able to optimize the model in an end-to-end manner.
2.2. Generative Retrieval
With the advancement of pre-trained language models, generative retrieval has emerged as an alternative to traditional document retrieval methods, establishing itself as a new paradigm. In contrast to traditional retrieval methods, generative retrieval leverages a single pre-trained generative model, such as T5 (Raffel et al., 2020) and BART (Lewis et al., 2020), to directly generate document identifier (docid) relevant to the query, enabling end-to-end optimization of the retrieval process. Additionally, this approach reduces reliance on external indexing, lowering the system’s demand for storage resources. Existing works have proposed various docid formats to represent the semantic information of documents, which can be categorized into two types: semantic docid and lexical docid. The semantic docid is represented as a sequence of numbers. These methods based on encoder-based PLM, like BERT (Devlin et al., 2019), to generate dense document embeddings. Afterward, techniques such as hierarchical k-means clustering (Tay et al., 2022; Wang et al., 2022; Sun et al., 2023) or product quantization (PQ) (Zhou et al., 2022) are applied to convert the document embeddings into sequential numeric form like “0-5-2”. This converted form serves as the semantic docid for the document, ensuring both efficient storage and effective retrieval based on the semantic structure, which is achieved by grouping semantically similar documents together. The lexical docid is a textual format that can well represent the semantic information of a document, and forms such as title (Cao et al., 2021), substrings (Bevilacqua et al., 2022), URL (Zhou et al., 2022), keywords (Wang et al., 2023; Zhang et al., 2023; Lee et al., 2023), and pseudo query (Tang et al., 2023; Li et al., 2023) are used to build the docid. Compared to methods that utilize semantic docid, methods based on lexical docid generate docid in a human-readable format, resulting in better interpretability. However, existing generative retrieval methods directly generate the docid for a user’s query, making it difficult to fully understand why the document is retrieved and how its rank is determined. To address these limitations, we propose a more reliable generative retrieval framework that not only generates relevant docid for the user’s query but also provides explanation.
Figure 2. Overview of HyPE framework. (1) HyPE constructs category paths using an external high-quality semantic hierarchy and employs LLM to select appropriate candidate paths for each document. (2) Then, HyPE links queries to the paths based on semantic relevance to construct path-augmented training set, and uses this to optimize the retrieval system. (3) During inference, HyPE employs path-aware reranking strategy to determine the final docid ranking by considering multiple paths.
3. Preliminaries
@@ 3.1. Task Formulation
(1) $\mathcal{M}^{\theta}(d\mid D)=\prod_{t=1}^{n}\mathcal{M}^{\theta}(d_{t}\mid D,d_{<t}),$
@@
(2) $\mathcal{M}^{\theta}(d\mid q)=\prod_{t=1}^{n}\mathcal{M}^{\theta}(d_{t}\mid q,d_{<t}),$
where $q$ is a query. In performing the aforementioned two tasks, it is important to 1) effectively represent the long document $D$ and 2) construct the docid $d$ that captures the overall semantic information of the document.
3.2. Document Representation and Identifier
3.2.1. Document representation
For the indexing task, each document is used as the input, making it crucial to define effective representations of the long document while preserving as much of its information as possible within the context length of the language model. The primary approaches to effectively representing documents are FirstP (Tay et al., 2022) and Document as Query (DaQ) (Wang et al., 2022). FirstP uses only the first $k$ tokens from the beginning of the document, while DaQ randomly extracts chunks from the entire document.
3.2.2. Document identifier
To ensure that docid effectively encodes semantic information of document, a variety of approaches have been proposed. As mentioned in section 2.2, docid can be broadly categorized into semantic docid and lexical docid. Semantic docid is a unique identifier, represented by either one number or a series of numbers. It is composed of document representations formed from encoder-based BERT, which are transformed into a compact numeric form. Lexical docid is a textual format designed to effectively convey the semantic content of a document. It can be constructed using various forms, such as the document’s title (Cao et al., 2021), substrings (Bevilacqua et al., 2022), keywords (Zhang et al., 2023; Lee et al., 2023; Wang et al., 2023), URL (Zhou et al., 2022), and pseudo query (Tang et al., 2023). Title and URL are used as docid directly from the dataset. Substrings are generated by the retrieval model using an FM index (Ferragina and Manzini, 2000), which creates specific n-grams within the document for retrieval. Keywords are extracted from the document using methods such as TF-IDF (Robertson and Walker, 1997), BM25 (Robertson and Zaragoza, 2009), or pre-trained language models (PLMs). Pseudo query is generated using query generation models such as docT5Query (Nogueira and Lin, 2020), and is then utilized as docid.
3.3. Optimization and Inference
3.3.1. Optimization via multi-task learning
@@
(3) $\max_{\theta}\sum_{(q,D,d)\in\mathcal{X}}\mathcal{M}^{\theta}(d\mid D)+\sum_{(q,D,d)\in\mathcal{X}}\mathcal{M}^{\theta}(d\mid q).$
3.3.2. Indexing with synthetic query
@@
In indexing task, documents are long and contain extensive information; on the contrary, in retrieval task, queries are relatively short and request specific information, highlighting a gap between the two tasks (Zhuang et al., 2023). To bridge this discrepancy, recent studies (Zhuang et al., 2023; Wang et al., 2022; Sun et al., 2023) have tried to integrate synthetic queries, generated by query generation (QG) models (Nogueira and Lin, 2020), into the training phase. The synthetic queries are helpful to improve the retrieval performance of generative retrieval models by effectively reducing the gap between queries and documents. Note that these synthetic queries are treated as alternative document representation, similar to FirstP and DaQ mentioned in 3.2.1, and are used as input for the indexing task (Pradeep et al., 2023; Zhuang et al., 2023; Sun et al., 2023).
3.3.3. Inference with constrained decoding
During inference, given an input query $q$, the model produces a top-$K$ ranked list of docids that have the largest likelihoods $\mathcal{M}^{\theta}(d\mid q)$. To ensure the generation of valid docids, the model employs constrained decoding, which mostly uses constrained beam search with a prefix trie (Cao et al., 2021).
4. Proposed Method
@@
In this section, we present a novel framework, Hierarchical category Path-Enhanced generative retrieval (HyPE), which enhances explainability by generating hierarchical category paths step-by-step before decoding docid. Our hierarchical category paths serve as pseudo-reasoning process, providing a reasonable and reliable explanation for why this document is retrieved. Furthermore, HyPE employs path-aware reranking strategy, which simultaneously considers multiple pseudo-reasoning paths to determine the documents to retrieve. The overall framework is illustrated in Figure 2.
4.1. Candidate Path Set Construction
The first step of our HyPE framework is to construct a set of candidate hierarchical category paths for each document. In this context, the hierarchical category paths have to satisfy the following criteria to ensure explainability:
  • Specificity: Are they capable of sufficiently explaining specific and detailed semantic information?
@@ 4.1. Candidate Path Set Construction
To achieve this, we first construct the high-quality backbone hierarchy for category paths in open-domain. Then, for each document, we (1) filter out category paths based on semantic similarity calculated by a pre-trained text encoder, and (2) select several category paths that comprehensively represent the content of the document while specifically addressing certain topics within the document by the help of outstanding reasoning capabilities of LLM.
4.1.1. Hierarchical category path collection.
@@
In the open-domain retrieval task, the category (or topic) hierarchy must encompass both a broad range of domain categories (i.e. width of tree) and sufficient semantic granularity (i.e. depth of tree) to ensure comprehensive and accurate retrieval system. To this end, we leverage Wikipedia’s category tree [^3] https://en.wikipedia.org/wiki/Special:CategoryTree as our backbone hierarchy of categories, setting the Main Topic classification category [^4] https://en.wikipedia.org/wiki/Category:Main_topic_classifications as the root node of the hierarchy. Considering the vast and complex nature of Wikipedia’s category tree, we limit the scraping process to a depth of four to construct our backbone hierarchy. Then, we linearize all the paths [^5] In this paper, for brevity, we use the term “path” to refer to a “category path”. within the hierarchy and convert them into a sequence of strings, thereby enabling more efficient processing and manipulation. The entire set of linearized category paths is denoted by $\mathcal{P}$. The statistics of collected hierarchical category paths are presented in Appendix A.1.
4.1.2. Candidate path set construction.
@@
The next step is to utilize the knowledge of LLM to assign appropriate category paths to each document within the corpus. However, due to the context length of LLM, it is impossible to input all possible paths within the category hierarchy (collected in Section 4.1.1);
thus, we first filter out path set for each document $D$ by leveraging a bi-encoder. [^6] We use Sentence-T5 (Ni et al., 2022) as the bi-encoder and cosine similarity as similarity score. The pre-candidate path set is obtained by
(4) $\hat{\mathcal{P}}_{D}=\operatorname*{argTop-}_{p\in\mathcal{P}}k\text{ sim}(E(D),E(p)),$
@@
where $E(\cdot)$ is the encoder, $\text{sim}(\cdot)$ is a similarity scoring function, and $k$ is the number of pre-candidate paths for each document. Then, we assign up to three paths that effectively represent the topics of the document by leveraging LLM few-shot prompting along with the document $D$ and its pre-candidate path set $\hat{\mathcal{P}}_{D}$  [^7] We use Llama-3-8B-Instruct as LLM..
(5) $\mathcal{P}_{D}=\text{LLM}({D},\hat{\mathcal{P}}_{D}).$
$\mathcal{P}_{D}$ refers to the final candidate path set of the document $D$. Our actual prompts used for this step can be found in Appendix A.4.
4.2. Optimization with Path-Augmented Dataset

### v2

@@ Abstract
Generative retrieval directly decode a document identifier (i.e., docid) in response to a query, making it impossible to provide users with explanations as an answer for “why is this document retrieved?”. To address this limitation, we propose Hierarchical Category Path-Enhanced Generative Retrieval (HyPE), which enhances explainability by first generating hierarchical category paths step-by-step then decoding docid. By leveraging hierarchical category paths which progress from broader to more specific semantic categories, HyPE can provide detailed explanation for its retrieval decision. For training, HyPE constructs category paths with external high-quality semantic hierarchy, leverages LLM to select appropriate candidate paths for each document, and optimizes the generative retrieval model with path-augmented dataset. During inference, HyPE utilizes path-aware ranking strategy to aggregate diverse topic information, allowing the most relevant documents to be prioritized in the final ranked list of docids. Our extensive experiments demonstrate that HyPE not only offers a high level of explainability but also improves the retrieval performance.
Why These Documents? Explainable Generative Retrieval with Hierarchical Category Paths
Sangam Lee1, Ryang Heo1, SeongKu Kang2, Susik Yoon2, Jinyoung Yeo1, Dongha Lee1thanks: Corresponding author 1Yonsei University   2Korea University {salee,ryang1119,jinyeo,donalee}@yonsei.ac.kr {seongkukang,susik}@korea.ac.kr
1 Introduction
Information retrieval (IR) systems are essential for helping users find proper information within vast amount of online information. A fundamental task of these systems is document retrieval, which focuses on searching for and ranking documents that are relevant to a given query from a large document corpus.
@@ 1 Introduction
Figure 1: Existing generative retrieval methods fail to explain why specific documents are retrieved, as they directly decode docid (Upper). In contrast, our HyPE provides clear explanations by generating query-related hierarchical category paths leading to the docid (Lower).
@@
Recently, generative retrieval has emerged as a new paradigm in document retrieval. It aims to directly generate document identifier (i.e., docid) for a given query by leveraging pre-trained generative models such as BART Lewis et al. (2020) and T5 Raffel et al. (2020). This paradigm enables end-to-end optimization of the retrieval process, allowing for fine-grained interaction between the input query and docid, and significantly reduces memory usage by leveraging the parametric memory of a single generative model.
@@
Even with these advantages, generative retrieval continues to face the challenge of determining how to construct docid that effectively represent documents. As the docid serves as a representation of the entire document, defining one that accurately encapsulates the document’s contents is both crucial and challenging. Existing works on generative retrieval have categorized docid into two types: semantic docid and lexical docid. A semantic docid represents each document as a series of numbers (e.g., 0-5-2), where each number indicates a cluster index assigned over its dense representation. This dense representation is encoded by a PLM-based encoder Devlin et al. (2019); Raffel et al. (2020) and clustered using methods such as hierarchical k-means Tay et al. (2022); Wang et al. (2022) or product quantization Zhou et al. (2022). On the other hand, lexical docid represents each document as human-readable text, such as titles Cao et al. (2021), keywords Zhang et al. (2023); Wang et al. (2023) and pseudo queries Tang et al. (2023).
@@
However, both existing approaches still lack explainability, which remains a significant limitation. For instance, in the upper part of Figure 1, two types of queries related to the same document “Dubai”, are presented. While the existing retrieval systems may return identifiers of relevant documents such as the lexical docid (i.e., Dubai) or semantic docid (i.e., 0-5-2), they fail to provide an explicit explanation that aligns with the different intention behind each query. Specifically, they do not clarify why a particular document is retrieved for a specific query and fail to answer the question, “why is this document retrieved?”. The lack of explainability in retrieval systems is a critical issue, as it can undermine the reliability of retrieved documents and make it more difficult for users to explore additional information related to a specific query Anand et al. (2022). To address this aforementioned limitation, our research aims to design a generative retrieval framework that can provide retrieved document with clear and reasonable explanations for a user’s query.
@@
In this work, we propose Hierarchical Category Path-Enhanced Generative Retrieval (HyPE), which enhances explainability by generating hierarchical category paths step-by-step before decoding docid. Motivated by structured document categorization systems, such as Wikipedia category tree or Microsoft Academic taxonomy Shen et al. (2018), HyPE utilizes hierarchical category paths as explanations, progressing from broad to specific semantic categories. In the lower part of Figure 1, when queries about document “Dubai” are given, HyPE uses category paths like “Government > Government by cities” or “Economy > Economy by cities” to explain why document “Dubai” is retrieved for each query. This approach 1) enables specific explanations for the document depending on the query by using hierarchical category paths that connect the query and the document, and 2) provides more reasonable and insightful explanation by reflecting the document’s semantic structure through a coarse-to-fine manner. Additionally, HyPE 3) can employ effective ranking of the retrieved results by leveraging multiple paths, which helps improve retrieval performance.
@@
Specifically, HyPE consists of the following three steps: 1) constructing category paths based on an external semantic hierarchy and selecting appropriate candidate paths for each document using Large Language Models (LLM), 2) building a path-augmented dataset with candidate paths, and 3) optimizing a model with the path-augmented dataset. During inference phase, HyPE conducts a pseudo-reasoning process [^1] We describe this term in Appendix A.2. by generating the hierarchical category path step-by-step to decode docid, allowing it to serve as an explanation which enhances explainability. Additionally, HyPE employs path-aware ranking strategy, which simultaneously considers multiple pseudo-reasoning paths for each query. This strategy helps build a more robust retrieval system by capturing the semantic information of multiple category paths, thereby improving overall retrieval performance.
@@
Our extensive experiments demonstrate that HyPE not only offers a high level of explainability but also improves the retrieval performance in the document retrieval task. Additionally, HyPE can be applied orthogonally to various docid types (e.g., title, keywords), making it a versatile framework that can be seamlessly integrated into different generative retrieval systems. For reproducibility, our codes are publicly available at the github repository: https://github.com/augustinLib/HyPE-Why-These-Documents.
We summarize our contributions as follows:
@@
  • We propose a new ranking strategy called path-aware ranking, which considers multiple category paths simultaneously to determine the final ranked list of docids.
@@
2 Preliminaries
@@ 2.1 Task Formulation
$\mathcal{M}^{\theta}(d\mid D)=\prod_{t=1}^{n}\mathcal{M}^{\theta}(d_{t}\mid D,d_{<t}),$ (1)
@@
$\mathcal{M}^{\theta}(d\mid q)=\prod_{t=1}^{n}\mathcal{M}^{\theta}(d_{t}\mid q,d_{<t}),$ (2)
where $q$ is a query. In performing the aforementioned two tasks, it is crucial to address two key aspects: 1) effectively represent the long document $D$ and 2) construct the docid $d$ that captures the overall semantic information of the document.
During inference, given an input query $q$, the model produces a top-$K$ ranked list of docids that have the largest likelihoods $\mathcal{M}^{\theta}(d\mid q)$. To ensure the generation of valid docids, the model employs constrained decoding, which mostly uses constrained beam search Cao et al. (2021).
Figure 2: Overview of HyPE framework. (1) HyPE constructs category paths using an external high-quality semantic hierarchy and employs LLM to select appropriate candidate paths for each document. (2) Then, HyPE links queries to the paths based on semantic relevance to construct path-augmented training set, and uses this to optimize the retrieval system. (3) During inference, HyPE employs path-aware ranking strategy to determine the final docid ranking by considering multiple paths.
2.2 Document Representation and Identifier
Document representation.
For the indexing task, each document is used as the input. This makes it crucial to define effective input representations of the long document while preserving as much of its information as possible within the context length of the language model. The primary approaches to effectively representing documents are FirstP Tay et al. (2022) and Document as Query (DaQ) Wang et al. (2022). FirstP uses only the first $k$ tokens from the beginning of the document, while DaQ randomly extracts chunks from the document.
Document identifier.
To ensure that docid effectively encodes semantic information of document, a variety of approaches have been proposed. Docid can be broadly categorized into semantic docid and lexical docid. Semantic docid represents each document as a series of numbers, where each number corresponds to a cluster index derived from the document’s dense representation. This dense representation is encoded by a PLM-based encoder Devlin et al. (2019) and mapped to discrete cluster indices using methods such as hierarchical k-means Tay et al. (2022); Wang et al. (2022) or product quantization Zhou et al. (2022). Lexical docid is a textual format designed to effectively convey the semantic content of a document. It can be constructed using various forms, such as the document’s title Cao et al. (2021), substrings Bevilacqua et al. (2022), keywords Zhang et al. (2023); Lee et al. (2023); Wang et al. (2023), URL Zhou et al. (2022), and pseudo query Tang et al. (2023). Title and URL are used as docid directly from the dataset. Substrings are generated by the retrieval model using an FM index Ferragina and Manzini (2000), which creates specific n-grams within the document for retrieval. Keywords are extracted from the document using methods such as TF-IDF Robertson and Walker (1997), BM25 Robertson and Zaragoza (2009), or pre-trained language models (PLMs). Pseudo query is generated using query generation models, such as docT5query Nogueira and Lin (2020), which is then utilized as the docid.
2.3 Optimization and Inference
Optimization via multi-task learning.
@@ 2.3 Optimization and Inference
$\max_{\theta}\sum_{(q,D,d)\in\mathcal{X}}\mathcal{M}^{\theta}(d\mid D)+\mathcal{M}^{\theta}(d\mid q)$ (3)
Indexing with synthetic query.
@@
In indexing task, documents are long and contain extensive information; however, in retrieval task, queries are relatively short and request specific information. To bridge this discrepancy, recent studies Zhuang et al. (2023); Wang et al. (2022); Sun et al. (2023) have tried to integrate synthetic queries, generated by query generation models Nogueira and Lin (2020), into the training phase. The synthetic queries improve the retrieval performance of generative retrieval models by effectively reducing the gap between queries and documents. Note that these synthetic queries are treated as alternative document representation, similar to FirstP and DaQ mentioned in 2.2, and are used as input for the indexing task Zhuang et al. (2023); Sun et al. (2023).
3 Proposed Method
@@
In this section, we present Hierarchical category Path-Enhanced generative retrieval (HyPE), which improves explainability by generating hierarchical category paths step-by-step before decoding docid. The overall framework is shown in Figure 2.
3.1 Candidate Path Set Construction
@@
The first step of our HyPE framework is to construct a set of candidate hierarchical category paths for each document. To ensure explainability, these paths should satisfy the following criteria: Semantic Hierarchy, Generalizability, and Specificity (see Appendix A.3 for details). To achieve this, we first construct the high-quality backbone hierarchy for category paths. Then, for each document, we (1) filter out category paths based on semantic similarity calculated by a pre-trained text encoder, and (2) select several category paths that comprehensively represent the content of the document while specifically addressing certain topics within the document by the help of reasoning capabilities of LLM.
Hierarchical category path collection.
@@
In the open-domain retrieval task, the category (or topic) hierarchy must encompass both a broad range of domain categories (i.e. width of tree) and sufficient semantic granularity (i.e. depth of tree) to ensure comprehensive and accurate retrieval system. To this end, we leverage Wikipedia’s category tree as our backbone hierarchy of categories, setting the Main Topic classification category as the root node of the hierarchy. This hierarchy is specifically designed to systematically categorize “real-world wikipedia documents”, which cover a wide range of domains and provide specific and detailed semantic information. Considering the vast and complex nature of Wikipedia’s category tree, we limit the scraping process to a depth of four to construct our backbone hierarchy. Then, we linearize all the paths within the hierarchy and convert them into a sequence of strings, thereby enabling more efficient processing and manipulation. The entire set of linearized category paths is denoted by $\mathcal{P}$. The statistics of collected hierarchical category paths are presented in Appendix A.3.
Candidate path set construction.
@@
Subsequently, we utilize the knowledge of LLM to assign appropriate category paths to each document within the corpus. However, due to the context length of LLM, it is impossible to input all possible paths within the category hierarchy (collected in Section 3.1). Thus, we first filter out path set for each document $D$ by leveraging a bi-encoder. The pre-candidate path set $\hat{\mathcal{P}}_{D}$ is obtained as follows:
$\hat{\mathcal{P}}_{D}=\operatorname*{argTop-}_{p\in\mathcal{P}}k\text{ sim}(E(D),E(p)),$ (4)
@@
where $E(\cdot)$ is the encoder, $\text{sim}(\cdot)$ is a cosine similarity, and $k$ is the number of pre-candidate paths for each document. Then, given the document $D$ and its pre-candidate path set $\hat{\mathcal{P}}_{D}$, we leverage LLM [^2] We use Llama-3-8B-Instruct Dubey et al. (2024) as LLM. to generate the final path set $\mathcal{P}_{D}$, selecting up to three paths that best represent the document.
3.2 Optimization with category path